OpenStudy (nonopro):

Subtract. 3/4 - 2/3 A.1/12 B.1/7 C.1/4 D.1/3

OpenStudy (anonymous):

wat do u think

OpenStudy (jadedry):

You'll have to make sure their denominators are equal before you subtract \[\frac{ 3 }{ 4} - \frac{ 2 }{ 3 } = \frac{ 9 }{ 12 } - \frac{ 8 }{ 12 }\] Multiply the numerators by the others denominators. Therefore the answer is: A

OpenStudy (alexandervonhumboldt2):

first make the denominators equal: 3/4-2/3=(3*3)/(4*3)-(2*4)/(3*4)=9/12-8/12=1/12